OpenStudy (anonymous):

y varies directly as to x. When x = 6 y = 2. Find y when x = 9.

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

xk=y 6k=2 k=2/3or 1/3 9(1/3) = y 9/3=y y=3

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

here is amother i will give u medal